# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v1

@@ 5 Experiments
In this section, we examine the performance of the two mLSTM variants mLSTMexp (mLSTM with exponential input gate) and mLSTMsig (mLSTM with sigmoid input gate). We compare two kernel algorithms: (1) limit_chunk: A kernel that is limited in chunk size $L$. (2) xl_chunk: Our TiledFlashLinearAttention (TFLA) kernels with unlimited chunk size. For details see Section 3.
We assess the performance of mLSTMsig compared to mLSTMexp in Section 5.1 and benchmark the runtime of our kernels against other baselines in Section 5.2. In App. E.1 we verify the numerical correctness of our kernels.
@@ 5.1 Language Modeling with mLSTM
We analyze the effect of the input gate bias initialization on training stability and performance of our mLSTM models in Appendix E.2. We observe in Figure 8 and 9, that initializing the input gate biases to -10 effectively mitigates large gradient norm spikes and variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 6), but also improves training stability, if initialized correctly.
@@
In Appendix E.2 we investigate the effect of the norm layer epsilon on language modeling performance for mLSTMexp. Our transfer behavior analysis in Figure 3 suggests, that there exists an interplay between norm layer epsilon and input gate bias initialization. We confirm this in our grid search in Figure 10 and find that the best performing configuration is the default epsilon ϵ=italic-$\epsilon=$1e-6 with input gate biases initialized to -10.
@@ 5.2 Kernel Benchmark
Finally, we compare the runtime of our mLSTM limit_chunk and TFLA xl_chunk kernels with kernel implementations of the state-of-the-art sequence modeling primitives FlashAttention (Dao, 2024; Shah et al., 2024), Mamba (Gu & Dao, 2024; Dao & Gu, 2024) and GLA (Yang et al., 2024b). In Appendix E.3 we compare with other kernels from the FlashLinearAttention library (Yang & Zhang, 2024). We run our benchmarks on NVIDIA H100 GPUs.
@@
We use the standard embedding dimension of 4096 for 7B parameter models and adapt the head dimensions per kernel accordingly. For example for FlashAttention we use 32 heads with head dim 128 and for the mLSTM we use 16 heads with head dim 256. Following the practice of Shah et al. (2024), we keep the number of tokens constant at 65,536 and vary sequence length and batch size accordingly. For further details see Appendix E.3.
Figure 4 shows the runtime benchmark results for forward pass only (left) and forward-backward pass (right). Our mLSTMexp TFLA xl_chunk kernels with two level sequence parallelism is about 25% faster than our limit_chunk kernels. Through targeted modifications of the input gate of the mLSTM we save computation and enable more efficient kernel implementations for the forward pass of mLSTMsig (see Sec. 4). This yields another speedup of over 30% for the forward pass of the mLSTMsig TFLA kernel over the mLSTMexp TFLA kernel. We perform additional runtime benchmarks for varying head dimensions and a more in-depth comparison to the FLA (Yang et al., 2024b) and LightningAttention2 (Qin et al., 2024a) kernels in Appendix E.3.
@@
The chunk size parameter $L$ balances the computation between the two levels of sequence parallelism (see Sec. 3.2). Smaller chunk sizes increase memory consumption because more chunks are materialized in memory, but they reduce the quadratic compute FLOPs in the parallel part. Larger chunk sizes have the opposite effect. They decrease memory consumption but increase quadratic compute FLOPs. In Figure 5 we measure this trade-off for our mLSTMsig TFLA xl_chunk kernels.
@@ 6 Related Work
TFLA builds on ideas from FlashAttention (Dao, 2024) and FlashLinearAttention (Yang et al., 2024b) and is designed for efficient mLSTM kernels (Beck et al., 2024), while being applicable also to other linear RNNs (Sun et al., 2023; Dao & Gu, 2024). We discuss this and other related works in Appendix A.
@@ 7 Conclusion and Future Work
Although we enhance training stability through careful gate initialization informed by our empirical transfer behavior analysis, future work could explore instabilities arising from numerical errors in kernel implementations in greater depth. Finally, the programming techniques and hardware features used to optimize FlashAttention (Shah et al., 2024) could also be leveraged by our TFLA algorithm. This makes us believe that TFLA has the potential to become a foundational primitive for future long-context language models.
@@ Acknowledgements
We thank Sebastian Böck, Richard Kurle, Patrick Blies, Antonio Orvieto, Maximilian Stadler, Thomas Schmied and Kajetan Schweighofer for helpful discussions and feedback.
@@ A.2 Other Related Work
Many different gating techniques for linear RNNs have been explored (Sun et al., 2023; Beck et al., 2024; Yang et al., 2024b; Gu & Dao, 2024; Dao & Gu, 2024; Qin et al., 2023, 2024b; Orvieto et al., 2023; Katsch, 2023). We propose mLSTMsig, a mLSTM variant with sigmoid input gate similar to the original LSTM (Hochreiter, 1991; Hochreiter & Schmidhuber, 1997) and analyze the transfer behavior, empirically.
@@
The mLSTM, which is the xLSTM with matrix memory, has already been adopted in several domains beyond language modeling. For example, Alkin et al. (2025) find that the mLSTM can serve as generic backbone for computer vision architectures. In the field of robotics, the mLSTM architecture has been explored as Large Recurrent Action Model, which leverages the efficient linear-time inference complexity of the mLSTM (Schmied et al., 2025), and as backbone for imitation learning policies (Jia et al., 2025). In the domain of biological and chemical sequences the mLSTM has been studied for generative modeling, representation and in-context learning (Schmidinger et al., 2025). Finally, the mLSTM has also been examined for in the domain time series forecasting (Alharthi & Mahmood, 2024), where also the sLSTM (Beck et al., 2024) has been applied (Kraus et al., 2024; Kong et al., 2025).
@@ B.2 Detailed Chunkwise-Parallel Formulation
In this section we provide more detailed formulas for the chunkwise-parallel formulation of the mLSTM from Section 2.2.
@@
Given the logarithmic forget gates $\bar{\bm{\mathrm{f}}}^{(k)}=\log\sigma(\tilde{\bm{\mathrm{f}}}^{(k)})\in\mathbb{R}^{L}$ and input gates $\bar{\bm{\mathrm{i}}}^{(k)}=\log\exp(\tilde{\bm{\mathrm{i}}}^{(k)})\in\mathbb{R}^{L}$, we can compute the chunkwise gates as
@@
$\bm{\mathrm{a}}_{k}$ $=\mathrm{rev\_cumsum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)+\bar{\bm{\mathrm{i}}}^{(k)}\;\in\mathbb{R}^{L},\;\text{with}\quad{\mathrm{a}}_{k,j}=\sum_{i=j+1}^{L}\,\bar{{\mathrm{f}}}^{(k)}_{i}+\bar{\mathrm{i}}^{(k)}_{j}\;\text{for}\;j=1,2,\dots,L.$ (38)
@@ D.2 Empirical Transfer Behavior Analysis of the mLSTM
In contrast, in Figure 7(a) and 7(b) we observe that incorporating a normalizer similar to mLSTMexp (excluding the max state) into mLSTMsig does not alter its transfer behavior.
The other two normalizer variants for mLSTMsig in Figure 7(c) and 7(d) show a clearly different transfer behavior and do not train successfully. Similarly, the variants in Figure 6(b) and 6(c) also fail to train successfully.
@@ E.2 Extended Language Modeling Experiments with mLSTM
Effect of Input Gate Bias Initialization (Figure 8 and 9).
In our transfer behavior analysis in Section 4.2 we find that there is a transition from suppressing the signal to passing the signal at negative input gate values of around -8 (see Figure 3). Since we initialize the weights of the gates $\bm{w}_{\{\mathrm{i},\mathrm{f}\}}$ to 0, the biases of the input and forget gates determine the actual position in the x-y plane in the beginning of training. Initially, with input gate biases initialized to 0, we observe a high gradient norm variance, which was more pronounced for mLSTMsig (see Figure 8(a) and 9(a)).
@@
Indeed, as we observe in Figure 8 and 9 initializing the input gate biases to -10 effectively mitigates gradient norm spikes and reduces high gradient norm variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 6), but also improves training stability, if initialized correctly.
@@
In our empirical transfer behavior analysis of the mLSTM in Section 4.2 and D.2 we find that the transfer behavior depends on the input and forget gate preactivations, as well as the normalization layer epsilon (see Figure 6(a) and 7(a)). Therefore, we perform a grid search over different normalization layer epsilons and input gate bias initializations for the mLSTM with exponential input gate with 160M parameters and 6 heads at context length 4096. We show the results in Figure 10.

### v2

@@ 5 Experiments
In this section, we examine the performance of the two mLSTM variants mLSTMexp (mLSTM with exponential input gate) and mLSTMsig (mLSTM with sigmoid input gate). We compare two kernel algorithms: (1) limit_chunk: A kernel that is limited in chunk size $L$. (2) xl_chunk: Our TiledFlashLinearAttention (TFLA) kernels with unlimited chunk size. For details see Section 3. We assess the performance of mLSTMsig compared to mLSTMexp in Section 5.1 and benchmark the runtime of our kernels against other baselines in Section 5.2. In App. E.1 we verify the numerical correctness of our kernels.
@@ 5.1 Language Modeling with mLSTM
We analyze the effect of the input gate bias initialization on training stability and performance of our mLSTM models in Appendix E.2. We observe in Figure 11 and 12, that initializing the input gate biases to -10 effectively mitigates large gradient norm spikes and variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 6), but also improves training stability, if initialized correctly.
@@
In Appendix E.2, we investigate the effect of the norm layer epsilon on language modeling performance for mLSTMexp. Our transfer behavior analysis in Figure 4 suggests, that there exists an interplay between norm layer epsilon and input gate bias initialization. We confirm this in our grid search in Figure 13 and find that the best performing configuration is the default epsilon ϵ=italic-$\epsilon=$1e-6 with input gate biases initialized to -10.
@@ 5.2 Kernel Benchmark
We compare the runtime of our mLSTM limit_chunk and TFLA xl_chunk kernels with kernel implementations of the state-of-the-art sequence modeling primitives FlashAttention (Dao, 2024; Shah et al., 2024), Mamba (Gu & Dao, 2024; Dao & Gu, 2024) and GLA (Yang et al., 2024b). In Appendix E.3 we compare with other kernels from the FlashLinearAttention library (Yang & Zhang, 2024). We run our benchmarks on NVIDIA H100 GPUs.
@@
We use the standard embedding dimension of 4096 for 7B parameter models and adapt the head dimensions per kernel accordingly. For example for FlashAttention we use 32 heads with head dim 128 and for the mLSTM we use 16 heads with head dim 256. Following the practice of Shah et al. (2024), we keep the number of tokens constant at 65,536 and vary sequence length and batch size accordingly. For further details see Appendix E.3. Figure 5 shows the runtime benchmark results for inference, i.e. forward pass only, (left) and for training, i.e. forward-backward pass (right). Our mLSTMexp TFLA xl_chunk kernels with two level sequence parallelism is about 25% faster than our limit_chunk kernels. Through targeted modifications of the input gate of the mLSTM we save computation and enable more efficient kernel implementations for the forward pass of mLSTMsig (see Sec. 4). This yields another speedup of over 30% for the forward pass of the mLSTMsig TFLA kernel over the mLSTMexp TFLA kernel.
In training, our TFLA kernels are faster than FlashAttention 3 for longer sequences and more than 2x faster than Mamba 2 kernels for all sequence lengths.
We perform additional runtime benchmarks for varying head dimensions and a more in-depth comparison to the FLA (Yang et al., 2024b) and LightningAttention2 (Qin et al., 2024a) kernels in Appendix E.3.
@@
The chunk size parameter $L$ balances the computation between the two levels of sequence parallelism (see Sec. 3.2). Smaller chunk sizes increase memory consumption, because more chunks are materialized in memory, but they reduce the quadratic compute FLOPs in the parallel part. Larger chunk sizes have the opposite effect. They decrease memory consumption, but increase quadratic compute FLOPs. In Figure 6, we measure this trade-off for our mLSTMsig TFLA xl_chunk kernels.
5.3 Theoretical Runtime and Arithmetic Intensity
In Figure 6, we empirically observe that there exists an optimal chunk size (between 128 and 256) at which the runtime is minimized. In Appendix G, we compute the theoretical runtime optimal chunksize and the arithmetic intensity of TFLA depending on the chunk size by taking the FLOPs, memory operations and hardware accelerator specification into account.
We find that the theoretical runtime optimum exceeds the empirically observed one (see Figure 20 in Appendix G.1), suggesting that our current kernel implementations may not yet fully exploit the available performance potential.
We analyze the effect of the chunk size parameter $L$ on runtime, FLOPs, and arithmetic intensity in detail in Appendix F and G summarize our findings as follows:
  1. 1.
    The chunk size $L$ mediates a trade-off between runtime and GPU memory usage. [Figure 6]
  2. 2.
    $L$ determines the total compute in FLOPs: $L=1$ matches the recurrent formulation, while $L=T$ matches the parallel one. [Figure 18]
  3. 3.
    There exists an optimal chunk size $L\in[1,T]$ that minimized the total FLOP count.
    [Equation (113), Figure 18, Figure 19]
  4. 4.
    Increasing $L$ raises the arithmetic intensity of TFLA kernels. [Equation (119), Figure 21]
  5. 5.
    The chunk size determines whether the kernel is memory-bound or compute-bound on a given hardware. [Figure 22, Figure 21]
  6. 6.
    FLOPs/s alone can be misleading; the optimal chunk size should be chosen based on total runtime.
    [Figure 23, Figure 20]
  7. 7.
    he runtime-optimal chunk size scales proportionally with the square root of the head dimension and the accelerator’s computational intensity.
    [Figure 24, Figure 25]
  8. 8.
    Newer hardware generations require larger chunk sizes to approach peak performance. [Figure 25, Figure 20]
@@ 6 Related Work
TFLA builds on concepts from FlashAttention (Dao, 2024) and Flash Linear Attention (Yang et al., 2024b), and is tailored for efficient mLSTM kernels (Beck et al., 2024). At the same time, it remains applicable to other linear RNNs such as RetNet, Mamba 2, and DeltaNet (Sun et al., 2023, 2024; Dao & Gu, 2024; Yang et al., 2024a). Additional discussion of these connections can be found in Appendix A.
@@ 7 Conclusion and Future Work
Although we enhance training stability through careful gate initialization informed by our empirical transfer behavior analysis, future work could explore instabilities arising from numerical errors in kernel implementations in greater depth. Finally, the programming techniques and hardware features used to optimize FlashAttention (Shah et al., 2024) could also be applied to our TFLA algorithm to approach peak performance on next-generation hardware, as suggested by our theoretical runtime analysis. This makes us believe that TFLA has the potential to become a foundational primitive for future long-context language models.
@@ Acknowledgements
We thank Sebastian Böck, Richard Kurle, Patrick Blies, Antonio Orvieto, Maximilian Stadler, Thomas Schmied, Kajetan Schweighofer and Sebastian Lehner for helpful discussions and feedback.
@@ A.2 Other Related Work
Many different gating techniques for linear RNNs have been explored (Sun et al., 2023; Beck et al., 2024; Yang et al., 2024b; Gu & Dao, 2024; Dao & Gu, 2024; Sun et al., 2024; Qin et al., 2024b; Orvieto et al., 2023; Katsch, 2023; Peng et al., 2024). We propose mLSTMsig, a variant of mLSTM with a sigmoid input gate similar to the original LSTM (Hochreiter, 1991; Hochreiter & Schmidhuber, 1997) and empirically analyze the transfer behavior.
@@
The mLSTM, which is the xLSTM with matrix memory, has already been adopted in several domains beyond language modeling. For example, Alkin et al. (2025) find that the mLSTM can serve as a generic backbone for computer vision architectures. In the field of robotics, the mLSTM architecture has been explored as a large recurring action model, which uses the efficient linear-time inference complexity of the mLSTM (Schmied et al., 2025), and as the backbone for imitation learning policies (Jia et al., 2025). In the domain of biological and chemical sequences, the mLSTM has been studied for generative modeling, representation, and in-context learning (Schmidinger et al., 2025). Finally, mLSTM has also been examined in the domain time series forecasting (Alharthi & Mahmood, 2024), where sLSTM (Beck et al., 2024) has also been applied (Kraus et al., 2024; Kong et al., 2025).
@@
A.3 Application of TFLA to other Linear RNNs
We have applied TFLA to linear RNNs with scalar headwise gates such as mLSTMexp and mLSTMsig. In this section, we show how TFLA could be applied to other Linear RNNs, but leave the implementation to future work.
Linear RNNs with Scalar Headwise Gates.
With minimal modifications, TFLA can be applied to other scalar headwise gated linear RNNs, such as, for example, Retention (Sun et al., 2023), Gated Retention (Sun et al., 2024), Simple GLA (Yang et al., 2024b) or Mamba 2 (Dao & Gu, 2024). In general, TFLA implementations of these linear RNNs can be obtained by modifying the forget and input gate parameterizations (of the TFLA mLSTMsig). Since neither of the aforementioned models has an input gate in the mLSTM sense, we fix the input gate of TFLA mLSTMsig to one (i.e. $\sigma(\tilde{\mathrm{i}}_{t})=1$). Then, for Simple GLA or Retention we keep the headwise scalar sigmoid forget gate $\sigma(\tilde{\mathrm{f}}_{t})$ or set it to a constant decay parameter $\sigma(\tilde{\mathrm{f}}_{t})=\gamma$.
To implement Gated Retention — which introduces a sigmoid forget gate modulated by a temperature parameter $\tau$ — we modify the activation function of the forget gate to incorporate this temperature term. Beyond coupling the forget gate parameter $\bm{A}$ with the keys $\bm{k}_{t}$ (represented by the parameter $\bm{B}$) through the step size parameter $\Delta$, Mamba 2 also adopts the linear attention structure from Equation (20). In this formulation, Mamba’s notation $\bm{C}$, $\bm{B}$, and $\bm{X}$ corresponds to our queries $\bm{Q}$, keys $\bm{K}$, and values $\bm{V}$, respectively.
Linear RNNs with Delta Rule.
Recently, Linear RNNs with non-diagonal state transition matrices have become popular due to their increased expressivity that enable them to solve synthetic state tracking tasks (Grazzi et al., 2025; Siems et al., 2025; Peng et al., 2025; Movahedi et al., 2025). One method to implement such non-diagonal linear RNNs efficiently is (Gated) DeltaNet (Yang et al., 2024c, a), which introduces a hardware efficient algorithm for linear transformers with the delta rule. The core of DeltaNet’s efficient implementation is the chunkwise-parallel formulation of the delta-rule, which is implemented using primitives from Flash Linear Attention.
In our notation, the chunkwise-parallel formulation of DeltaNet is given by the inter-chunk recurrence
$\underset{(d_{qk}\times d_{hv})}{\bm{C}_{k}}=\underset{(d_{qk}\times d_{hv})}{\bm{C}_{k-1}}+\underset{(d_{qk}\times L_{kv})}{{\bm{K}^{(k)}}^{\top}}\ \underbrace{\left(\underset{(L_{kv}\times d_{hv})}{\bm{U}^{(k)}}-\underset{(L_{kv}\times d_{qk})}{\bm{W}^{(k)}}\ \underset{(d_{qk}\times d_{hv})}{\bm{C}_{k-1}}\right)}_{\bm{V}^{*(k)}},$ (27)
and the combination between inter-chunk and intra-chunk contribution
$\underset{(L_{hq}\times d_{hv})}{\bm{\mathrm{H}}^{(k)}}=\underset{(L_{hq}\times d_{qk})}{\bm{Q}^{(k)}}\ \underset{(d_{qk}\times d_{hv})}{\bm{C}_{k-1}}+\left(\underset{(L_{hq}\times d_{qk})}{\bm{Q}^{(k)}}\ \underset{(d_{qk}\times L_{kv})}{{\bm{K}^{(k)}}^{\top}}\right)\ \underbrace{\left(\underset{(L_{kv}\times d_{hv})}{\bm{U}^{(k)}}-\underset{(L_{kv}\times d_{qk})}{\bm{W}^{(k)}}\ \underset{(d_{qk}\times d_{hv})}{\bm{C}_{k-1}}\right)}_{\bm{V}^{*(k)}},$ (28)
where we omit the mask $\bm{\mathrm{M}}$ applied to $\bm{Q}\bm{K}^{\top}$ of DeltaNet for clarity. We refer to Yang et al. (2024c) for the definition of the matrices $\bm{U}^{(k)}$ and $\bm{W}^{(k)}$.
By defining the new values $\bm{V}^{*(k)}=\bm{U}^{(k)}-\bm{W}^{(k)}\bm{C}_{k-1}$ we can recover the core formula of TFLA for the intra-chunk parallel forward pass (see Equation (20)) to which we can apply our TFLA tiling strategy from Section 3.2. For computing the matrices $\bm{U}^{(k)}$, $\bm{W}^{(k)}$ and the new values $\bm{V}^{*(k)}$ one could follow the same strategy as in DeltaNet, where $\bm{U}^{(k)}$ and $\bm{W}^{(k)}$ are computed in a separate kernel before the recurrent kernel. Then, the new values $\bm{V}^{*(k)}$ are computed and stored in HBM in the recurrent kernel together with the memory states $\bm{C}_{k}$.
We conclude that TFLA can be applied to any linear RNN that either follows or can be reformulated into the simplified, chunkwise-parallel form of Equation (20).
@@ B.2 Detailed Chunkwise-Parallel Formulation
In this section, we provide more detailed formulas for the chunkwise-parallel formulation of the mLSTM from Section 2.2.
@@
Given the logarithmic forget gates $\bar{\bm{\mathrm{f}}}^{(k)}=\log\sigma(\tilde{\bm{\mathrm{f}}}^{(k)})\in\mathbb{R}^{L}$ and input gates $\bar{\bm{\mathrm{i}}}^{(k)}=\log\exp(\tilde{\bm{\mathrm{i}}}^{(k)})\in\mathbb{R}^{L}$, we can compute the logarithmic chunkwise gates as
@@
$\bm{\mathrm{a}}_{k}$ $=\mathrm{rev\_cumsum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\texttt{[1:]}\right)+\bar{\bm{\mathrm{i}}}^{(k)}\;\in\mathbb{R}^{L},\;\text{with}\quad{\mathrm{a}}_{k,j}=\sum_{i=j+1}^{L}\,\bar{{\mathrm{f}}}^{(k)}_{i}+\bar{\mathrm{i}}^{(k)}_{j}\;\text{for}\;j=1,2,\dots,L\ ,$ (40)
where [1:] denotes (in numpy notation) that the first index is excluded as it is done in the sum notation of Equation (40). Additionally, in Figure 8 we illustrate the chunkwise gate computation and show a PyTorch code snippet for computing the chunkwise gates in Figure 7.
1def compute_chunkwise_log_gates_vecB_vecA(
2 vecI: torch.Tensor, # (B, NH, S)
3 vecF: torch.Tensor, # (B, NH, S)
4 chunk_size: int,
5):
6 B, NH, S = vecI.shape
7 assert S %
8 _device = vecI.device
9 NC = S // chunk_size
10 L = chunk_size
11
12 # compute vecB
13 vecF_logsig = logsigmoid(vecF.to(dtype=torch.float32))
14 vecF_logsig_chunked = rearrange(vecF_logsig, "b nh (nc l) -> b nh nc l", nc=NC, l=L)
15 vecB = vecF_logsig_chunked.cumsum(dim=-1)
16
17 # compute vecA
18 vecI_chunked = rearrange(vecI, "b nh (nc l) -> b nh nc l", nc=NC, l=L)
19 # unstable vecA computation:
20 # vecA = (vecB[..., -1, None] - vecB) + vecI # (B, NH, NC, L)
21 # stable vecA computation:
22 vecA = (
23 torch.cat(
24 [
25 vecF_logsig_chunked[..., 1:].flip(-1).cumsum(-1).flip(-1),
26 torch.zeros((B, NH, NC, 1), device=_device, dtype=torch.float32),
27 ],
28 dim=-1,
29 )
30 + vecI_chunked
31 ) # (B, NH, NC, L)
32 return vecB, vecA
Figure 7: PyTorch function to compute the logarithmic chunkwise gates for mLSTMexp.
Figure 8: Illustration of the chunkwise gate computation.
@@ D.2 Empirical Transfer Behavior Analysis of the mLSTM
In contrast, in Figure 10(a) and 10(b) we observe that incorporating a normalizer similar to mLSTMexp (excluding the max state) into mLSTMsig does not alter its transfer behavior.
The other two normalizer variants for mLSTMsig in Figure 10(c) and 10(d) show a clearly different transfer behavior and do not train successfully. Similarly, the variants in Figure 9(b) and 9(c) also fail to train successfully.
@@ E.2 Extended Language Modeling Experiments with mLSTM
Effect of Input Gate Bias Initialization (Figure 11 and 12).
In our transfer behavior analysis in Section 4.2 we find that there is a transition from suppressing the signal to passing the signal at negative input gate values of around -8 (see Figure 4). Since we initialize the weights of the gates $\bm{w}_{\{\mathrm{i},\mathrm{f}\}}$ to 0, the biases of the input and forget gates determine the actual position in the x-y plane in the beginning of training. Initially, with input gate biases initialized to 0, we observe a high gradient norm variance, which was more pronounced for mLSTMsig (see Figure 11(a) and 12(a)).
@@
Indeed, as we observe in Figure 11 and 12 initializing the input gate biases to -10 effectively mitigates gradient norm spikes and reduces high gradient norm variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 6), but also improves training stability, if initialized correctly.
@@
In our empirical transfer behavior analysis of the mLSTM in Section 4.2 and D.2 we find that the transfer behavior depends on the input and forget gate preactivations, as well as the normalization layer epsilon (see Figure 9(a) and 10(a)). Therefore, we perform a grid search over different normalization layer epsilons and input gate bias initializations for the mLSTM with exponential input gate with 160M parameters and 6 heads at context length 4096. We show the results in Figure 13.